OpenStudy (anonymous):

Charlie has 8 bills in his wallet. Five of the bills are $10 bills, and the other bills are $5 bills. What is the total value of the 8 bills in Charlie's wallet? The total value of the 8 bills is $ .

OpenStudy (acxbox22):

5 x 10 + 3 x 5= 50 + 15= $65